# High-dimensional dynamics in low-dimensional networks

### v1

@@ Results
    be the singular value decomposition of $W$ with $U$ and $V$ being $N\times r$ orthonormal matrices. We define the $r\times r$ recurrent alignment matrix [3, 24],
    $P=U^{T}V.$
    which measures the alignment between the left and right singular vectors of $W_{0}$, i.e., the columns of $U$ and $V$. High-dimensional dynamics can only emerge when no singular value of $P$ is small, i.e., when $\sigma_{P}=\mathcal{O}(1)$ for all singular values of $P$.
@@
To better understand this condition on $P$, we consider simulations of a model that is widely used in computational neuroscience to model a recurrently interacting network of neurons [16, 25, 2, 26],
@@
We first consider the special case that $W_{0}$ is normal, so ${\bf{v}}_{k}=\pm{\bf{u}}_{k}$. In this case, $P$ is diagonal with $|P_{kk}|=1$ (Figure 2b) so all singular values of $P$ are $\sigma_{P}=1$ and the network exhibits high-dimensional dynamics (Figure 2c). Since $W_{0}$ is normal, its eigenvalues are real and equal in magnitude to the singular values, $|\lambda_{W_{0}}|=\sigma_{W_{0}}$. Since $\sigma_{W_{0}}\gg 1$, stability of the dynamics requires that $\lambda_{W_{0}}=-\sigma_{W_{0}}$ and therefore
$V$ $=\left[\begin{array}[]{cc}-{\bf{u}}_{2}&-{\bf{u}}_{2}\end{array}\right].$
Dynamics on the column space of $U$ are strongly suppressed due to the large, negative eigenvalues.
Figure 2: Conditions for high-dimensional dynamics in a network with rank-two structure. a) Singular values of $W$ demonstrate an effective low-rank structure. b) The alignment matrix, $P$, when $W_{0}$ is normal. c) The variance explained by each principal component of the inputs, ${\bf{x}}(t)$, and network response, ${\bf{z}}(t)$, demonstrates high-dimensional dynamics. d,e) Same as b,c except $W_{0}$ is EP, but not normal. f,g) Same as a,b except the first left singular vector is orthogonal to all right singular vectors, so $W_{0}$ is not EP. h,i) Same as a,b except the left and right singular vectors of $W_{0}$ are only partially aligned (Eq. (4) with $c=0.5$).
However, low-rank suppression does not require that $W_{0}$ is normal (${\bf{v}}_{k}=\pm{\bf{u}}_{k}$). Instead, it is sufficient that $U$ and $V$ share a column space, $\textrm{col}(U)=\textrm{col}(V)$. Low-rank matrices, $W_{0}=U\Sigma V^{T}$, for which $U$ and $V$ share a column space are called “equal projector” (EP) matrices [27]. All normal matrices are EP, but an EP matrix is not necessarily normal. An example of a non-normal, EP matrix with rank 2 is given by
@@
The resulting matrix $W_{0}=U\Sigma V^{T}$ is highly non-normal because ${\bf{u}}_{1}\cdot{\bf{v}}_{1}={\bf{u}}_{2}\cdot{\bf{v}}_{2}=0$, $P$ is zero along the diagonal (Figure 2d), and $|\lambda_{W_{0}}|\neq\sigma_{W_{0}}$. Regardless, all singular values of $P$ are $\sigma_{P}=1$, and therefore the network produces low-rank suppression and high-dimensional dynamics (Figure 2e).
To obtain an example of a network that produces low-dimensional dynamics, we take
@@
where ${\bf{v}}_{\perp}$ is orthogonal to ${\bf{u}}_{1}$ and ${\bf{u}}_{2}$. In this case, $P$ has one singular value at $\sigma_{P}=0$ and another at $\sigma_{P}=1$ (Figure 2f). Because of the singular value at zero, the network produces low-dimensional dynamics in which one principal component captures an outsized proportion of the variability (Figure 2g).
Finally, we consider the case in which left and right singular vectors are not perfectly aligned, but have some non-vanishing overlap, for example
@@
where $10<c<$ 1. In this case, $P$ has a singular value at $\sigma_{P}=\sqrt{c}$ and another at $\sigma_{P}=1$ (Figure 2h). As long as $c$ is not asymptotically close to zero, the network produces low-rank suppression and high-dimensional dynamics (Figure 2i).
In summary, the class of effectively low-rank matrices, $W$, that produce low-dimensional dynamics in the presence of high-dimensional inputs is the very specific class of matrices for which the alignment matrix, $P=U^{T}V$, of the low-rank part is singular or nearly singular in the sense that it has a small or zero singular value.
It is common in theoretical work to consider models in which the entries of $U$ and $V$ are random and independent with zero mean. In this case, ${\bf{u}}_{j}$ and ${\bf{v}}_{k}$ are nearly orthogonal so $\sigma_{P}\ll 1$ and the network produces low-dimensional dynamics (Supplementary Figure S.2). However, many networks arising in nature and applications do not have purely random structure. In the following sections, we show that many naturally arising network structures satisfy our conditions for high-dimensional dynamics.
@@
So far, we considered example networks unbiased weights, $E[W_{jk}]=0$, which is common in many modeling studies, but many networks in nature have weights with non-zero mean. Biased weights can produce low-rank structure. As a simple example, consider a random network with independent weights satisfying
@@
If $m$ and $s$ scale similarly to each other, then the largest singular value of $W$ is near $\sigma_{1}=|m|N$ while the next-largest singular value scales like $\sigma_{2}\sim\mathcal{O}(s\sqrt{N})$, implying an effective rank-one structure because $\sigma_{1}/\sigma_{2}\sim\mathcal{O}(\sqrt{N})$. The dominant rank-one part of such a matrix has constant entries, so it is normal and the network exhibits low-rank suppression and high-dimensional responses to high-dimensional perturbations (Supplementary Figure S.3).
@@
where $W^{a,b}$ is a sub-matrix quantifying connections from population $b$ to population $a$. In general, each sub-matrix can have a different, non-zero mean and variance,
$E[W_{jk}]=m_{ab}\;\textrm{ and }\;\textrm{std}(W_{jk})=s_{ab}.$
@@
If we assume that each population has $\mathcal{O}(N)$ members and that $m_{ab}$ scale similarly to $s_{ab}$ for large $N$, then $W$ has up to $n$ dominant singular values that scale like $\mathcal{O}(m_{ab}N)$ and the remaining singular values scale like $\mathcal{O}(\sigma_{ab}\sqrt{N})$. Hence, large modular networks with biased weights naturally produce low-rank structure [1].
@@
Networks of this form can be decomposed as $W=W_{0}+W_{1}$ where $W_{0}=E[W]$ is the element-wise expectation of $W_{0}$, which is constant within each block, so $W_{0}$ has rank at most $n$. In general, $W_{0}$ is not a normal matrix, but it is an EP matrix because $\textrm{col}(U)$ and $\textrm{col}(V)$ are each spanned by the $n$ indicator vectors of the $n$ populations. Therefore, modular networks with biased weights exhibit low-rank suppression and high-dimensional dynamics in response to high-dimensional perturbations.
As a specific example of a modular network, we consider a model of a local neuronal network in the cerebral cortex. Cortical neurons obey Dale’s Law: All outgoing connection weights from a particular neuron have the same sign, positive for excitatory neurons and negative for inhibitory neurons, and mean connection weights also depend on the postsynaptic neuron type [29, 30]. These properties produce a modular structure in which the columns of the adjacency matrix corresponding to excitatory neurons are non-negative, while the columns corresponding to inhibitory neurons are non-positive (Figure 3a). We consider the case in which $m_{ab}$ and $s_{ab}$ scale like $1/\sqrt{N}$, consistent with experiments [31] and theoretical work [32, 33, 34, 35, 3, 36].
@@
This network has an effective rank-two structure (Figure 3b): Two singular values scale like $\sqrt{N}$ while the others are $\mathcal{O}(1)$ and the low-rank part of the connectivity matrix is EP (Figure 3c), so our theory predicts that the network exhibits low-rank suppression and high-dimensional responses to high-dimensional inputs. The column space of $U$ and $V$ consist of all vectors that are uniform within each population,
$\textrm{col}(U)=\textrm{col}(V)=\{[a\;\cdots\;a\;b\;\cdots\;b]^{T}\,|\,a,b\in{\mathbb{R}}\}.$
@@
Therefore, perturbations that are uniform within each population are aligned to the low-rank part and suppressed relative to random perturbations (Figure 3d,e) consistent with recordings from monkey motor cortex [36] and previous theoretical work on balanced network models [37, 38, 39, 36].
As predicted, simulations also show high-dimensional responses to high-dimensional perturbations (Supplementary Figure S.4), consistent with observations that neural responses in primate visual cortex are high-dimensional when visual stimuli are high-dimensional [13].
@@
Averaging over the excitatory and inhibitory populations represents a projection onto $\textrm{col}(U)$. As a result, the cancellation mechanism illustrated in Figure 1g manifests as a tight balance between mean excitatory (positive) and inhibitory (negative) input to neurons (Figure 3f), a phenomenon that is widely observed in neural recordings [6, 7, 8, 9, 10, 11, 12] and widely studied in computational models [32, 33, 39, 3]. Therefore, the widely studied phenomenon of excitatory-inhibitory balance emerges as a consequence of low-rank suppression.
@@
Many networks in nature exhibit connection strength that depends on the distance between nodes in physical or other spaces, resulting in an effective low-rank structure [1, 34]. As a simple example, we considered a model in which each node is assigned a location, $\theta\in[0,1)$, and connection strength decays like a Guassian function of the distance, $d\theta$, between nodes so that only nearby nodes are more strongly connected (Figure 4a). Connectivity was also perturbed by a random component, $W_{1}$, as above. This connectivity structure is effectively low-rank with the first several singular values proportional to the Fourier coefficients of the Gaussian connectivity kernel while the remaining singular values, inherited from $W_{1}$, are much smaller (Figure 4b). The low-rank part is normal, with left- and right singular vectors equal to the Fourier basis vectors (Supplementary Figure S.5), so the network satisfies our conditions for low-rank suppression and high-dimensional responses.
Perturbations that are smooth in space are aligned to the low-rank part of the connectivity matrix because they are formed by sums of low frequency spatial Fourier modes. Therefore, due to low-rank suppression, network responses to spatially disordered perturbations are amplified relative to smooth perturbations (Figure 4c-e; compare tick labels on vertical axes). The network also exhibits high-dimensional responses to high-dimensional perturbations (Supplementary Figure S.6), as predicted.
@@
We next consider a real epidemiological network, specifically a network of high school social contacts [40], which was used in a recent theoretical study of low-rank networks [1]. In that study, the authors considered quenched mean-field reduction of the susceptible-infected-susceptible epidemiological model,
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ W{\bf{z}}$ (5)
where $g>0$, $\gamma>0$ and $\circ$ denotes element-wise multiplication. In this model, each ${\bf{z}}_{j}(t)$ models the probability that an individual is infected. Here, $W$ is the proximity matrix of 637 high school students, indicating whether each pair of students were in proximity of each other during a specific school week [40]. This matrix is effectively low-rank in the sense that it has a small number of dominant singular values (Supplementary Figure S.10a). In [1], it was shown that the dynamics, ${\bf{z}}(t)$, generated by Eq. (5) on this network are effectively low-dimensional. However, Eq. (5) is completely self contained without any internal noise or external perturbations. Noise and perturbations arise in epidemiological dynamics through interactions with individuals from outside of the modeled network (for example, parents and siblings outside of the high school) and through the natural stochasticity of infection.
Since the proximity matrix is necessarily symmetric and therefore normal, our analysis predicts that the network will produce low-rank suppression and high-dimensional responses to high-dimensional perturbations. To test this prediction, we modified the model by adding an external forcing term,
@@
Simulations of these dynamics indeed demonstrate low-rank suppression (Figure 5a and Supplementary Figure S.10b) and high-dimensional dynamics (Figure 5b) in contrast to the model from [1]. These results highlight the importance of accounting for external perturbations and internal noise when studying the dimensionality of epidemiological dynamics. Moreover, the results imply that epidemiological networks are, perhaps counterintuitively, more sensitive to random perturbations than to perturbations aligned with the network’s low-rank structure.

### v2

@@ Results
where $\Sigma$ is a diagonal, $r\times r$ matrix of singular values ($\Sigma_{kk}=\sigma_{k}\gg 1$) while $U=[{\bf{u}}_{1}\cdots{\bf{u}}_{r}]$ and $V=[{\bf{v}}_{1}\cdots{\bf{v}}_{r}]$ are $N\times r$ orthonormal matrices with columns ${\bf{u}}_{k}$ and ${\bf{v}}_{k}$ that define the left- and right-singular vectors of $W_{0}$.
We additionally assume that $W_{1}$ is a random matrix that is independent of $U$ and $V$. In other words, $W$ is a low-rank matrix perturbed by noise. This is a stronger assumption than necessary ($W_{1}$ need not be fully independent from $W_{0}$ for our results to hold) but it simplifies our analysis.
[…]
Low-rank suppression occurs under Assumptions 1 and 2.
Our first result is that Assumptions 1 and 2 are sufficient for low-rank suppression (as observed in Figures 1d and 2d). Specifically,
Claim 1.
Consider the dynamics defined by Eq. (6) under Assumptions 1 and 2. There exists an input ${\bf{x}}_{a}$ with $\|{\bf{x}}_{a}\|=\mathcal{O}(1)$ such that the network’s steady-state response to ${\bf{x}}_{a}$ is asymptotically larger than the $\mathcal{O}(1)$ response to a random input, ${\bf{x}}_{\textrm{r}}$ (where $\|{\bf{x}}_{\textrm{r}}\|=\mathcal{O}(1)$).
To show why Claim 1 is true, first choose a $k\in\{1,\ldots,r\}$ and define
${\bf{x}}_{a}={\bf{u}}_{k}+\frac{W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}.$ (8)
Note that $\|{\bf{x}}_{a}\|=1+o(1)$ since $\sigma_{k}\gg 1$, $\|{\bf{u}}_{k}\|=\|{\bf{v}}_{k}\|=1$, and the maximum singular value of $W_{1}$ is $\mathcal{O}(1)$. Now note that the steady-state response to the input ${\bf{x}}_{a}$ is given by
${\bf{z}}_{a}=-\frac{{\bf{v}}_{k}}{\sigma_{k}}.$
This can be checked by substituting the equations for ${\bf{x}}_{a}$ and ${\bf{z}}_{a}$ into the steady-state equation, ${\bf{z}}=W{\bf{z}}+{\bf{x}}$. Since $\|{\bf{v}}_{k}\|=1$ and $\sigma_{k}\gg 1$, we have that $\|{\bf{z}}_{a}\|\ll 1$. Since $\sigma_{k}\gg 1$, the network strongly suppresses the input ${\bf{x}}_{a}$ from Eq. (8). In summary, low-rank suppression is a generic phenomenon in linearly stable, strongly low-rank networks, regardless of whether connectivity is symmetric or normal. This explains why low-rank suppression was observed in both examples from Figure 1 and Figure 2.
Note from Eq. (8) that ${\bf{x}}_{a}$ is closely aligned to ${\bf{u}}_{k}$ since $\sigma_{k}\gg 1$. Hence, the suppressed input, ${\bf{x}}_{a}$, is nearly (but not exactly) parallel to the left singular vector, ${\bf{u}}_{k}$. Also note that the response, ${\bf{z}}_{a}$, is perfectly aligned to ${\bf{v}}_{k}$. Finally, note that there are $r$ different choices for ${\bf{x}}_{a}$ (one for each $k\in\{1,\ldots,r\}$), all of which are nearly orthogonal to each other.
Interestingly, the input-response directions along which suppression occurs are reversed from the input-response directions along which amplification occurs in a feedforward network with the same connectivity (for example, the network in Supplementary Figure S.1 for which ${\bf{z}}=W{\bf{x}}$ in the steady-state). More specifically, in a feedforward network with connectivity $W$, inputs aligned to the dominant right singular vectors (${\bf{v}}_{k}$) produce amplified responses in the direction of the corresponding left singular vectors (${\bf{u}}_{k}$). This is a left-right reversal and a suppression-amplification reversal of the behavior for recurrent networks, for which inputs aligned to ${\bf{u}}_{k}$ produce suppressed responses in the direction of ${\bf{v}}_{k}$.
In summary, recurrent networks with strongly low-rank structure satisfying Assumptions 1 and 2 admit a small number of input directions along which inputs are strongly suppressed. We call this form of suppression low-rank suppression. In later sections, we will explore consequences of low-rank suppression in common low-rank network structures. Before discussing the dimensionality of dynamics, we first need to make an additional assumption about the dimensionality of the input.
Assumption 3: External inputs are high-dimensional and slowly varying.
To begin discussing the dimensionality of ${\bf{z}}(t)$, we need a model of ${\bf{x}}(t)$ that allows us to quantify dimensionality in a concrete way. To this end, we assume that ${\bf{x}}(t)$ is a stationary ergodic stochastic process in which case ${\bf{z}}(t)$, as defined in Eq. (6), is also a stationary ergodic process whenever dynamics are stable [31]. This allows quantify the dimensionality of ${\bf{x}}(t)$ and ${\bf{z}}(t)$ in terms of their principal components or, equivalently, by the eigenvalues of their stationary covariance matrices.
We assume that ${\bf{x}}(t)$ is high-dimensional in the sense that the singular values of the covariance matrix of ${\bf{x}}(t)$ (i.e., the variance explained by the principle components of ${\bf{x}}(t)$) do not have a small number of dominant terms (as in Figure 1d; see Supplementary Text S.2 for more mathematical details).
The assumption that ${\bf{x}}(t)$ is high-dimensional means that we are focused on the question of whether the network intrinsically generates low-dimensional dynamics. When ${\bf{x}}(t)$ is low-dimensional, low-dimensional dynamics in ${\bf{z}}(t)$ can be inherited from ${\bf{x}}(t)$ instead of being generated intrinsically (see Supplementary Figure S.2 for an example).
We additionally assume that ${\bf{x}}(t)$ varies in time more slowly than the intrinsic timescale, $\tau$, of the network interactions. To make this assumption more precise, note that whenever ${\bf{x}}(t)={\bf{x}}$ is constant, solutions, ${\bf{z}}(t)$, converge to the steady-state given by ${\bf{z}}=[I-W]^{-1}{\bf{x}}$. We assume that ${\bf{x}}(t)$ varies sufficiently slowly that ${\bf{z}}(t)$ is approximated by the quasi-steady state approximation,
${\bf{z}}(t)\approx[I-W]^{-1}{\bf{x}}(t).$
In Supplementary Text S.2, we give a more mathematically detailed definition of this assumption. In Supplementary Text S.3, we also provide a preliminary generalization of our analysis to the case where ${\bf{x}}(t)$ is not assumed to vary slowly, in which case the dimensionality of transient dynamics can be analyzed in the Laplace domain.
In simulations, we take each ${\bf{x}}_{j}(t)$ to be a smooth, i.i.d. Gaussian process with correlation time $\tau_{x}=10$. Our assumption of high-dimensionality is satisfied because each ${\bf{x}}_{j}(t)$ is independent. Our assumption that ${\bf{x}}(t)$ varies slowly is satisfied because $\tau_{x}$ is much larger than the network’s intrinsic timescale, $\tau=1$.
Conditions for low-dimensional dynamics.
In Supplementary Text S.2, we prove that (under Assumptions 1, 2, and 3) the dimensionality of dynamics depends on the recurrent alignment matrix [3] of $W_{0}$, defined by
$P=V^{T}U$
This $r\times r$ matrix measures the alignment between the left and right singular vectors of $W_{0}$. Specifically, $P_{jk}={\bf{v}}_{j}\cdot{\bf{u}}_{k}$ measures the alignment between the $j$th right singular vector and the $k$th left singular vector of $W_{0}$. This matrix is also sometimes referred to as the “overlap matrix” [11]. Our main result from Supplementary Text S.2 can be stated as follows:
Claim 2.
Consider the dynamics defined by Eq. (4) under Assumptions 1–3. If the dynamics of ${\bf{z}}(t)$ are low-dimensional then $P=V^{T}U$ has at least one asymptotically small singular value.
Here, “low-dimensional dynamics” is interpreted to mean that the covariance matrix of ${\bf{z}}(t)$ has an $\mathcal{O}(1)$ number of dominant eigenvalues, equivalently, ${\bf{z}}(t)$ has a small number of dominant principal component directions. Note that this result is equivalent to the following:
Claim 2.
Consider the dynamics defined by Eq. (4) under Assumptions 1–3. If all singular values of $P=V^{T}U$ are $\mathcal{O}(1)$ then the dynamics of ${\bf{z}}(t)$ are high-dimensional.
Recall that all singular values of $P$ are bounded by $1$, so the condition that all singular values of $P$ are $\mathcal{O}(1)$ simply says that no singular values of $P$ are asymptotically small, hence the two versions of Claim 2 are logical contrapositives.
Figure 3: Conditions for high-dimensional dynamics in a network with rank-two structure. a) Singular values of $W$ demonstrate an effective low-rank structure. b) The alignment matrix, $P$, when $W_{0}$ is normal. c) The variance explained by each principal component of the inputs, ${\bf{x}}(t)$, and network response, ${\bf{z}}(t)$, demonstrates high-dimensional dynamics. d,e) Same as b,c except $W_{0}$ is EP, but non-normal. f,g) Same as a,b except the first left singular vector is orthogonal to all right singular vectors, so $W_{0}$ is not EP. h,i) Same as a,b except the left and right singular vectors of $W_{0}$ are only partially aligned (Eq. (10) with $c=0.5$).
@@
To better understand the conditions on $P$ in Claim 2, we consider simulations of a widely used model from computational neuroscience [27, 32, 2, 33],
@@
We first consider a case in which $W_{0}$ is normal by setting
$V$ $=\left[\begin{array}[]{cc}-{\bf{u}}_{1}&-{\bf{u}}_{2}\end{array}\right].$
This is an extension of the example from Figure 1 to rank $r=2$. In this case, $P$ is diagonal with $P_{kk}=-1$ (Figure 3b) so all singular values of $P$ are $\sigma_{P}=1$ and the network produces high-dimensional dynamics (Figure 3c).
In this example, $W_{0}$ is both normal and symmetric. Results are similar when $W_{0}$ is normal and asymmetric (obtained, for example, by applying a rotation to $V$ in $r=2$ dimensions) except the conclusion $|P_{kk}|=1$ only holds when the non-zero eigenvalues of $W_{0}$ are real, which is necessarily the case when the $\sigma_{k}$ are distinct.
However, high-dimensional dynamics do not require that $W_{0}$ is normal. Instead, it is sufficient that $U$ and $V$ share a column space, $\textrm{col}(U)=\textrm{col}(V)$. Low-rank matrices, $W_{0}=U\Sigma V^{T}$, for which $U$ and $V$ share a column space are called “equal projector” (EP) matrices [34]. All normal matrices are EP, but an EP matrix is not necessarily normal. If $W_{0}$ is an EP matrix, then $\sigma_{P}=1$ for all singular values of $P$, so the condition in Claim 2 is satisfied. An example of a non-normal, EP matrix $W_{0}$ with rank 2 is given by taking
@@
The resulting matrix $W_{0}=U\Sigma V^{T}$ is highly non-normal because ${\bf{u}}_{1}\cdot{\bf{v}}_{1}={\bf{u}}_{2}\cdot{\bf{v}}_{2}=0$, $P$ is zero along the diagonal (Figure 3d), and $|\lambda_{W_{0}}|\neq\sigma_{W_{0}}$. Regardless, all singular values of $P$ are $\sigma_{P}=1$, so the network produces high-dimensional dynamics (Figure 3e).
To obtain an example of a network that produces low-dimensional dynamics (see also the example in Figure 2), we set
@@
where $\bf{v}$$_{\perp}$ is orthogonal to ${\bf{u}}_{1}$ and ${\bf{u}}_{2}$, so $P$ has one singular value at $\sigma_{P}=0$ and another at $\sigma_{P}=1$ (Figure 3f). Because of the singular value at zero, the condition in Claim 3 is not met, and the network produces low-dimensional dynamics in which one principal component captures an outsized proportion of the variability (Figure 3g).
Finally, we consider the case in which left and right singular vectors are not perfectly aligned, but have some non-vanishing overlap,
@@
where $0<c<1$. In this case, $W_{0}$ is not EP, but $P$ has a singular value at $\sigma_{P}=\sqrt{c}$ and the other at $\sigma_{P}=1$ (Figure 3h). As long as $c$ is not close to zero, the condition in Claim 2 is satisfied, and the network produces high-dimensional dynamics (Figure 3i).
It is common in theoretical work to consider models in which the entries of $U$ and $V$ are random and independent with zero mean. In this case, ${\bf{u}}_{j}$ and ${\bf{v}}_{k}$ are nearly orthogonal so $\sigma_{P}\ll 1$ and the network produces low-dimensional dynamics (Supplementary Figure S.3; compare to Figure 2). However, many networks arising in nature do not have purely random structure. In the following sections, we show that many naturally arising network structures satisfy our conditions for high-dimensional dynamics.
@@
So far, we considered networks with unbiased weights, $E[W_{jk}]=0$, which is common in modeling studies, but many networks in nature have weights with non-zero mean. Biased weights can produce low-rank structure. As a simple example, consider a random network with independent weights satisfying
@@
If $m$ and $s$ scale similarly, then the largest singular value of $W$ is near $\sigma_{1}=|m|N$ while the next-largest singular value scales like $\sigma_{2}=\mathcal{O}(s\sqrt{N})$, implying an effective rank-one structure when $N$ is large. The dominant rank-one part has constant entries, so it is normal and the network exhibits low-rank suppression and high-dimensional dynamics (Supplementary Figure S.6; see Appendix B in [35] for related results).
@@
where $W^{a,b}$ is a $N_{a}\times N_{b}$ sub-matrix quantifying connections from population $b$ to population $a$. In general, each sub-matrix can have a different, non-zero mean and variance,
$E\left[W^{a,b}_{jk}\right]=m_{ab}\;\textrm{ and }\;\textrm{std}\left(W^{a,b}_{jk}\right)=s_{ab}.$
@@
If we assume that each population has $N_{a}\sim\mathcal{O}(N)$ members and that the $m_{ab}$ scale similarly to the $s_{ab}$, then $W$ has up to $n$ dominant singular values that scale like $\mathcal{O}(m_{ab}N)$ and the remaining singular values scale like $\mathcal{O}(s_{ab}\sqrt{N})$. Hence, when $s_{ab}\sim m_{ab}$, large modular networks with biased weights naturally produce a low-rank structure in the sense that a small number of singular values are asymptotically larger than the rest [1]. If additionally $m_{ab}\gg\mathcal{O}(1/N)$, then the dominant singular values are asymptotically large, so the network is strongly low-rank in the sense defined in Assumption 2.
@@
Networks of this form can be decomposed as $W=W_{0}+W_{1}$ where $W_{0}=E[W]$ is constant within each block, so $W_{0}$ has rank at most $n$. In general, $W_{0}$ is not a normal matrix, but it is an EP matrix because $\textrm{col}(U)$ and $\textrm{col}(V)$ are each spanned by the indicator vectors of the $n$ populations,
$\textrm{col}(U)=\textrm{col}(V)=\textrm{span}\left\{{\bf{1}}_{1},{\bf{1}}_{2},\ldots,{\bf{1}}_{n}\right\}$
where each ${\bf{1}}_{k}$ is an $N$-dimensional indicator vector with entries defined by
${\bf{1}}_{k}]_{j}=\begin{cases}1&\textrm{if index $j$ is in population $k$}\\ 0&\textrm{otherwise}\end{cases}$
Therefore, modular networks with biased weights exhibit low-rank suppression and high-dimensional dynamics.
[…]
@@
Together, these biologically justified assumptions imply that the network has a strongly low rank structure, as defined in Assumption 2, with rank $r=2$ (Figure 4b). Specifically, Two singular values scale like $\sqrt{N}$ while the others are $\mathcal{O}(1)$. Moreover, the low-rank part of the connectivity matrix is EP (Figure 4c). Hence, the network produces high-dimensional dynamics and low-rank suppression despite the fact that $W$ is low-dimensional and non-normal. The column space of $U$ and $V$ consist of all vectors that are uniform within each population,
$\textrm{col}(U)=\textrm{col}(V)$ $=\textrm{span}\{{\bf{1}}_{e},{\bf{1}}_{i}\}$
$=\{[a\;\cdots\;a\;b\;\cdots\;b]^{T}\,|\,a,b\in{\mathbb{R}}\}$
@@
where ${\bf{1}}_{e}=\{1,2,\ldots,N_{e}\}$ and ${\bf{1}}_{i}=\{N_{e}+1,N_{e}+2,\ldots,N\}$. Therefore, perturbations that are uniform within each population are aligned to the low-rank part and suppressed relative to random perturbations (Figure 4d,e). In other words, perturbations that stimulate all excitatory neurons equally and all inhibitory neurons equally are suppressed relative to perturbations that are inhomogeneous within one or both populations, consistent with previous work on balanced network models of cortical circuit responses [43, 23, 22, 24]. As predicted, simulations also show high-dimensional responses to high-dimensional inputs (Supplementary Figure S.7), consistent with observations that neural responses in monkey visual cortex are high-dimensional when visual stimuli are high-dimensional [16].
@@
Averaging over the excitatory and inhibitory populations represents a projection onto $\textrm{col}(U)=\textrm{col}(V)$. As a result, the cancellation mechanism illustrated in Figure 1h manifests as a tight balance between mean excitatory (positive) and inhibitory (negative) input to neurons (Figure 4f), a phenomenon that is widely observed in neural recordings [44, 45, 46, 47, 48, 49, 50] and widely studied in computational models [21, 40, 24, 3]. Hence, the widely studied theory of balanced network models can be interpreted as a special case of the low-rank suppression and low-rank cancellation studied here, specific to a particular modular network structure.
@@
Many networks in nature exhibit connection strength that depends smoothly on the distance between nodes in physical or other spaces, resulting in an effective low-rank structure [1, 41]. As a simple example, we consider a model in which each node is assigned a spatial location, $\theta\in[0,1)$, and connection strength decays like a Guassian function of the distance, $d\theta$, between nodes so that only nearby nodes are strongly connected (Figure 5a). Connectivity is also perturbed by a random component, $W_{1}$, as above. This connectivity structure is effectively low-rank (Figure 5b) and the low-rank part is normal, with left- and right singular vectors spanned by the low-frequency Fourier basis vectors (Supplementary Figure S.8). Therefore, networks of this form satisfy our conditions for low-rank suppression and high-dimensional dynamics (Supplementary Figure S.9).
Perturbations that are smooth in space are aligned to the low-rank part of the connectivity matrix because they are formed by sums of low-frequency spatial Fourier modes. Hence, perhaps surprisingly, strongly low-rank networks with spatial structure are more sensitive to spatially disordered perturbations than to spatially smooth perturbations (Figure 5c-e; compare tick labels on vertical axes).
@@
We next consider a real epidemiological network, specifically a network of high school social contacts [51], which was used in recent theoretical work on low-rank network dynamics [1]. In that work, the authors considered a quenched mean-field reduction of the susceptible-infected-susceptible model,
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ W{\bf{z}}$ (11)
where $g>0$, $\gamma>0$, and $\circ$ denotes element-wise multiplication. Each ${\bf{z}}_{j}(t)$ models the probability that an individual is infected. Here, $W$ is the proximity matrix of 637 high school students, indicating whether each pair of students were in proximity of each other during a specific school week [51]. This matrix is effectively low-rank in the sense that it has a small number of dominant singular values (Supplementary Figure S.10a). In [1], it was shown that the dynamics generated by Eq. (11) on this network are effectively low-dimensional.
@@
However, Eq. (11) is completely self-contained without any perturbations. Perturbations arise in epidemiological dynamics through interactions with individuals from outside of the modeled network (for example, parents and siblings outside of the school) and through the natural stochasticity of infection.
Since the proximity matrix is necessarily symmetric and therefore normal, our analysis predicts that the network produces low-rank suppression and high-dimensional responses to high-dimensional perturbations. To test this prediction, we modified the model by adding an external forcing term,
@@
Simulations of these dynamics indeed demonstrated low-rank suppression (Figure 6a and Supplementary Figure S.10b) and high-dimensional dynamics (Figure 6b) in contrast to the model from [1]. These results highlight the importance of accounting for external perturbations and internal noise when studying the dimensionality of epidemiological dynamics. Moreover, the results imply that epidemiological networks are, perhaps counterintuitively, more sensitive to random perturbations than to perturbations aligned with the network’s low-rank structure.